Tuesday, September 21, 2010

BELL RINGER

A{Q, R, S, T, U, V} B{T, U, V, W, X, Y, Z}

1. How many elements are in {A}? n{A} = 6
2. How many elements are in {B}? n{B} = 7
3. how many elements are in {AuB}? n{AuB} = 10 because T, U, V are in both sets, so they are only counted once
4. How many elements are in {AnB}? n{AuB} = 3 because T, U, V are in both sets
5. Are these two sets independent or dependent? These sets are dependent because they overlap over elements T, U, V

AGENDA
1. Bellringer
2. Homework Review
3. Independent probabilities

OBJECTIVE
Students will learn to identify independent probabilities

HOMEWORK
Begin review set A on pg. 490. Remember to ask questions tomorrow for the test on Thursday.

QUESTION
P(G)=0.3, P(D)=0.7 and P(GuD)= 1, are G and D dependent or independent?


I know, horrible question... Sorry this was so late. My computer has a virus or something so i would get half done and it would crash, leaving me to start over... Hooray for attempt number 11! Yay!

hey

guys I just got my username

Monday, September 20, 2010

September 20, 2010

Bell Ringer
1: If the probability of an Event E is 40%, what is the probability of not E? (E')
2: If the same event E takes place 3 times, what is the probability of the result being
E E' E' ??

1. Probability of E is 40% = 4/10 ....which leaves the probability of E' to be 6/10.
6+4= 10

2. k so its really hard to do a tree diagram, but that would be the best step so, time to go to plan B which iiissss...just explaining it.... ok so
  • first we have to calculate the probability of E which is 4/10
  • then, we know that the probability of E' is 6/10
  • again, we know that the probability of E' is 6/10
  • So now we can MULTIPLYYYYY
4/10 * 6/10 * 6/10 = 8/125

So the answer for the probability of E E' E' is 18/125

Agenda
I. Bell Ringer
II. Theoretical vs. Experimental Probability
III. Independent Events (in other words... h.w..we got tricked! =] )

Homework!!
pg. 490 #1-5

OBJECTIVE
Students will be able to differentiate the difference between experimental and theoretical Probability.


Question of the day....drum roll please!!!!

1.The probability that Blazer will spend the day with a Princess is 80%. What is the probability that Blazer won't spend the day with the Princess?

2. Over three days, what is the probability that Blazer will spend the first two days with the princess and not on the third day? (P P P' )







Thursday, September 16, 2010

September 16, 2010

Agenda

1) Bell Ringer
2) Homework review
3) Independent Events
4) Review for Quiz

Objective: Students should be able to identify dependent and independent events and calculate probability accordingly.

Bell ringer( we went over it in class)

Biology

Physics

Chemistry

Total

Male

40

16

35

91

Female

15

24

20

59

Total

55

40

55

150


1) P(M)= 91/150
2) P(M) ∪ P(C)= 91+20/150 = 111/150
3) P(P) ∩ P(M)= 16/91

Homework Review
During class, we went over question 6 from page 483.

You are supposed to draw a venn diagram to solve the problem. (I'm so sorry- I tried many different ways to copy my venn diagram from microsoft word but it won't work.)

The venn diagram has two circles
1st circle is labeled Mu for mumps and the 2nd circle is labeled as Me for Measles.
The part where the two circles overlap(in the middle) is labeled as Me and Me because it contains both
So, here are the values
Mu= 2
Me=12
Mu and Me(middle)= 12
Outside the circle but inside the box (neither mu or me)= 4
All these values add upto 30.

a) P(Mu)= 14/30
b) P(Mu) - P(Mu
∩ Me)
24/30 -14/30 = 2/30
(you don't have to reduce it because IB doesn't expect you to reduce it on the exams)
c) P(Mu'
∪ Me')= 4/30
d) P(Mu
∩ Me)/ P(Mu)= 12/14

Tomorrow's Quiz
- probability question including chart, tree diagram, and venn diagram.

September 23, 2010: Chapter test for Chap. 14.

Practice question:


Out of 30 kids, 20 love strawberry and chocolate, 24 love chocolate and 23 love strawberry.

a) How many kids love neither of the flavors mentioned above?
b) What is the probability that a kid loves chocolate ONLY?
c) What is the probability that a kid loves chocolate given he loves strawberry?
d) What would you pick? Strawberry or Chocolate? :)

Hey everyone :)
I'm so sorry for not posting the venn diagram :\ I really tried to for the past hour but it didn't work. If you have any questions about today's blog, please let me know. I will be more than glad to help you out! I hope the practice question was easy. Do not forget- we have a quiz tomorrow!!!
Have a blessed day guys.

Lillian Sangha

Wednesday, September 15, 2010

Math Studies II

AGENDA:
I) Bell Ringer
II) Venn Diagrams to evaluate probability
III) Math Studies Project Timeline



I) Bell Ringer

80 students were asked what type of television program they watched the previous night.

35 - watched sports (S)
50 - watched drama (D)
42 - watched news (N)
10 - watched all 3 programs
7 - watched sports and news only
12 - watched news and drama only
14 - watched sports and drama only

*Draw a Venn diagram of the above problem.

SOLUTION:



* To solve the problem it would be good to follow the letters:

a) 10 goes into center of all 3 circles as everyone watched it.

b) 7 goes between sport and news

c) 12 goes between news and drama

d) 14 goes between sports and drama

e) To find "e", we have to add all the numbers that are included in "sport circle" which is red color, and that includes ( a)10 + b)7 + d)14 + e)? = 35, as 35 students watched sports program.

10 + 7 + 14 + E = 35
31 + E = 35
E = 4
------------------------------------------------
*We do the same procedure with drama:
To find "F" we need to add: a)10 + c)12 + d)14 + f)? = 50

10 + 12 + 14 + F = 50
36 + F = 50
F = 14
------------------------------------------------
*And News:

To find " G" we need to add: a)10 + b)7 + c)12 + g)? = 42

10 + 7 + 12 + G = 42
29 + G = 42
G = 13
--------------------------------------------------------------------
*After completing those three subproblems we will obtain this:


However, this isn't the and of the problem, because we have to find h. To find it, we first have to add all the numbers, and they should add up to 80. a) 10 + b)7 + c)12 + d)14 + e)4 + f)14 + g)13 + h)? = 80
10 + 7 + 12 + 14 + 4 + 14 + 13 + H = 80
74 + H = 80
H = 6

*Here is the final solution:



Some Classwork:


1) The probability that the student is boy in P.E. = 37/70
2) The probability that boy GIVEN English = 10/18
3) The probability Math GIVEN boy = 12/42
4) The probability Girl GIVEN P.E. = 15/35


OBJECTIVE:
Students will be able to use Venn diagrams to solve probability problems.


HOMEWORK:

IB Math Studies Book Page 483,#4-7.


QUESTIONS:

1) In IB classes there are 52 students. 30
students love Math, 10 ART,and 15 TOK.
3 students like Art and Math
2 students like Math Art and TOK
1 student like Art and TOK
5 student like Math and TOK.

Find out how many students like Math alone, Art and TOK and how many students, if any, don't like any of the subject.

2) Which group are You in?


Tuesday, September 14, 2010

Math Math Math.. (I forgot the name of the lesson)

BELLRINGER

Create a decision tree for:

Probability it will snow is .75
If it snows, the probability that I will ski is .9
Probability I will ski if it doesn't snow is .2
Find the Probability that I will ski


ANSWER: 29/40


SOLUTION:



Ski .9
/
.75 Snow
/ \
/ .1 No Ski
/
.25 No Snow /-- .2 Ski
/ \/
/
.8 No Ski


To find the probability of skiing you:

1. Multiply the .75 probability of snow X the .9 of skiing if there's snow which equals .675 and put into fraction form is 27/40.
2. Multiply the .25 probability of no snow X .2 of skiing without snow and that equals 2/40.
3. Multiply the .1 probability of not skiing with snow X .75 of snow which equals 3/40.
4. Multiply one last time the .25 probability of no snow X not skiing without snow and that equals 8/40.
5. Then you add 27/40 + 2/40 and that equals 29/40, which is the probability you will ski.



AGENDA

I. Bell Ringer
II. Review Homework
III. Probability w/ Venn Diagram
IV. Schedule for Math IA




Objective:
Student will be able to use Venn Diagram to illustrate real world situations and find a probability.




Question:
Bob lives in a house. Everytime it rains heavily his basement floods and he has to take out buckets of water. Probability it will rain heavy is .4. Probability it won't rain is .6. Probability he won't take out buckets of water if there's no rain is .6.
What is the probability he will take out buckets of water?



Ahh. I know this is a little late.. I barely got home half an hour ago. Easiest question I could come up with.. lol.
Have fun. The tree looked better before I posted it.

Monday, September 13, 2010

September 13, 2010

BELLRINGER (which was confusing)

Create a tree diagram given 2 archers Li and Yuka. Li has a 3/4 probability of hitting. Yuka has a 4/5 probability of hitting. Show all possible outcomes for each making two shots at target.

SOLUTION:
H STANDS FOR HIT
M STANDS FOR MISS

LI
/ \
3/4 H 1/4 M
/ \ / \
4/5 H 1/5 M 4/5 H 1/5 M
/ \ / \ / \ / \
3/4 H 1/4 M 3/4 H 1/4 M 3/4 H 1/4 M 3/4 H 1/4 M
/ \ / \ / \ / \ / \ / \ / \ / \
4/5H 1/5M 4/5H 1/5M 4/5H 1/5M 4/5H 1/5M 4/5H 1/5M 4/5H 1/5M 4/5H 1/5M 4/5H 1/5M
Possible Outcomes: (Start left, work your way down)

When Li First Hits it
1. 3*4*3*4 = 144/400 (H,H,H,H)
2. 3*4*3*1 = 36/400 (H,H,H,M)
3. 3*4*1*4 = 48/400 (H,H,M,H)
4. 3*4*1*1 = 12/400 (H,H,M,M)
5. 3*1*3*4 = 36/400 (H,M,H,H)
6. 3*1*3*1 = 9/400 (H,M,H,M)
7. 3*1*1*4 = 12/400 (H,M,M,H)
8. 3*1*1*1 = 3/400 (H,M,M,M)

When Li First Misses
9. 1*4*3*4 = 48/400 (M,H,H,H)
10. 1*4*3*1 = 12/400 (M,H,H,M)
11. 1*4*1*4 = 16/400 (M,H,M,H)
12. 1*4*1*1 = 4/400 (M,H,M,M)
13. 1*1*3*4 = 12/400 (M,M,H,H)
14. 1*1*3*1 = 3/400 (M,M,H,M)
15. 1*1*1*4 = 4/400 (M,M,M,H)
16. 1*1*1*1 = 1/400 (M,M,M,M)




AGENDA

I. Bell Ringer
II. Sample Space
III. Independent vs Dependent Events

OBJECTIVE
Students will be able to use tree diagrams to show sample space and define independent and Dependent Events

HOMEWORK
IB Math Studies Book Page 488 #s 9, 10 , 11

Question to Answer (if you want, remember you have to answer 2 each week)

Lucia goes to lunch. She can either choose white or chocolate milk. And for the snack she can choose between chocolate and sugar cookies. List ALL of her possible outcomes.
Just a fun question, which combination would you pick and why?

:) . This took forever, i hope you guys understand and enjoy. see you guys i have to go now.